How many grams are needed of ammonium hydroxide if it is reacted with bismuth(III) sulfate to give you 16 grams of bismuth hydroxide, the other product is ammonium sulfate. Also how many grams of bismuth(III) sulfate are needed?
1
Expert's answer
2020-04-10T13:36:40-0400

The reaction can be represented as following:

6NH4OH + Bi2(SO4)3 --> 2Bi(OH)3 + 3(NH4)2SO4

From the reaction:

m(NH4OH) / [6 × Mr(NH4OH)] = m(Bi(OH)3) / [2 × Mr(Bi(OH)3)] = m(Bi2(SO4)3) / Mr(Bi2(SO4)3)

From here:

m(NH4OH) = [m(Bi(OH)3) × 6 × Mr(NH4OH)] / [2 × Mr(Bi(OH)3)] = [16g × 6 × 35] / [2 × 260 g/mol] = 6.46 g

m(Bi2(SO4)3) = [m(Bi(OH)3) × Mr(Bi2(SO4)3)] / [2 × Mr(Bi(OH)3)] = [16 g × 706 g/mol] / [2 × 260 g/mol] = 21.72 g


Answer: 6.46 g of NH4OH and 21.72 g of Bi2(SO4)3.